# Flash Inference: Near Linear Time Inference for Long Convolution Sequence Models and Beyond

### v1

@@ 3.1.2 Performing LCSM inference via fast relaxed polynomial interpolation
We can now use Algorithm 1 to get an efficient inference algorithm for LCSMs, but to do so we need to be careful about data-dependency. First, for brevity, in our previous notation of activations, disregard the dependency on input $x$ that is, use $a_{\ell,i}$ instead of $(a_{\ell}(x))_{i}$ since we are only going to work with one input $x$ (the one we are generating). Furthermore, denote the intermediate mixer computation by $b\in\mathbb{R}^{M\times L\times D}$ as follows:
$b_{\ell,i}\triangleq\mathrm{mixer}_{\ell}(a_{\ell-1})_{i}=\sum_{k=0}^{i}{a_{\ell-1,k}\odot\rho_{\ell,i-k}}$ (4)
$a_{\ell,i}\triangleq\mathrm{block}_{\ell}(b_{\ell,i})$ (5)
for any $1\leq\ell\leq M$, where $a_{0,i}$ is the embedding of $i$th token. Whenever one has generated the tokens $x_{[1,i]}$ up to position $i$ or, equivalently, $a_{0,[1,i]}$, the whole of $a_{[0,M]\times[1,i-1]}$ will have been computed (since $a_{M,[1,i-1]}$ is needed to sample $x_{[1,i-1]}$). Caching these up is the equivalent of a KV-cache in transformers and is simply the natural way to not repeat work when autoregressively sampling. To generate the next token $x_{i+1}=a_{0,i+1}$, we want to fill in the values of $a_{[1,M],i}$. To do so, as per Eq. 4, we need to compute $b_{\ell,i}$ which is practically, for each level $\ell$, and within each dimension $1\leq c\leq D$, a relaxed interpolation between $a_{\ell-1,:,c}$ (playing the role of $y$) and $\rho_{\ell,:,c}$. Further applying a $\mathrm{block}_{\ell}$ to $b_{\ell}$, one gets $a_{\ell}$. Thus, we can pipeline several instances of Algorithm 1 sequentially across layers.
The resulting algorithm is illustrated in Algorithm 2. The outer loop (line 3) iterates through positions $i$: at each iteration, one computes the activations at position $i$ across all layers, does some (limited) eager work and samples next token. To do so, the inner loop (line 5) iterates through the several mixer-layers. It first accounts for the red cells in the diagram, that is, the direct dependency on previous layer’s activations at position $i$ (line 7), thus finalizing $b_{\ell,i}$. After computing the activation at position $i$ (line 8), it also accounts for the contribution of the gray tile that has just been unlocked (line 10) at the current layer. We reuse the notation from previous section for $\tau$ to be the same as before but within each of the $D$ dimensions - that is:
@@
0:  filter $\rho\in\mathbb{R}^{M\times L\times D}$, $\mathrm{block}_{[1,M]}$, first token $a_{0,1}$ and $\mathrm{sampler}$
1:  Output: All activations $a_{0},\ldots,a_{M}\in\mathbb{R}^{L\times D}$ obtained by autoregressively sampling
@@
3:  for $i\leftarrow 1$ to $L-1$ do
4:     $U\leftarrow$ maximum power of $2$ that divides $i$ the side of the $i$th gray tile
@@
6:        # account for the contribution of $a_{\ell-1,i}$ to $b_{\ell,i}$ red cell
7:        $b_{\ell,i}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=a_{\ell-1,i}\odot\rho_{\ell,0}$
8:        $a_{\ell,i}=\mathrm{block}_{\ell}(b_{\ell,i})$
9:        # account for the contribution of $a_{\ell-1,[i-U+1,i]}$ to $b_{\ell,[i+1,i+U]}$ gray tile
10:        $b_{\ell,[i+1,i+U]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=\tau(a_{\ell-1},[i-U+1,i],\rho_{\ell},[i+1,i+U])$
@@
The overall time complexity of Algorithm 2 is $O(MDL\log^{2}L)$ for the mixer-part plus $LM$ $\mathrm{block}$ calls, to generate $L=2^{P}$ tokens. Furthermore, for each $10\leq q\leq P-$ 1, there are $MD2^{P-1-q}$ calls to $\tau$ on ranges of length $2^{q}$. The overall amount of memory it takes to store the activations is $O(MLD)$ (as is the case for the naive lazy approach) and the peak memory usage does not change the asymptotics.
@@ 3.2 Across-layer Parallelization
One important feature of the proposed method is that it allows for higher parallelization across layers. In particular, the gray lines can be taken out of the inner loop and performed in parallel across all layers at once, just after the loop (Algorithm 3) - this is because all their inputs and outputs are disjoint. However, the red lines need to be performed sequentially since $a_{\ell,i}$ is a function of $b_{\ell,i}$ which is a function of $a_{\ell-1,i}$ for every $\ell$.
@@
0:  filter $\rho\in\mathbb{R}^{M\times L\times D}$, $\mathrm{block}_{[1,M]}$, first token $a_{0,1}$ and $\mathrm{sampler}$
1:  Output: All activations $a_{0},\ldots,a_{M}\in\mathbb{R}^{L\times D}$ obtained by autoregressively sampling
@@
3:  for $i\leftarrow 1$ to $L-1$ do
4:     $U\leftarrow$ maximum power of $2$ that divides $i$ the side of the $i$th gray tile
@@
6:        # account for the contribution of $a_{\ell-1,i}$ to $b_{\ell,i}$ red cell
7:        $b_{\ell,i}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=a_{\ell-1,i}\odot\rho_{\ell,0}$
8:        $a_{\ell,i}=\mathrm{block}_{\ell}(b_{\ell,i})$
@@
10:     # account for the contribution of $a_{\cdot,[i-U+1,i]}$ to $b_{\cdot,[i+1,i+U]}$ gray tile(s)
@@
12:         $b_{\ell,[i+1,i+U]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=\tau(a_{\ell-1},[i-U+1,i],\rho_{\ell},[i+1,i+U])$
@@ 3.3 Memory Considerations
As follows from Proposition 1, the overall size of the inputs to (and outputs of) $\tau$ in the proposed algorithm is only $O(MDL\log L)$ that is, we access on average activation values at $\log L$ positions per iteration as opposed to the naive implementations (such as lazy and eager approaches) that access an average of $\Omega(L)$ positions. This directly translates to data movement improvements by a factor of up to $L/\log L$.
A simple optimization of static memory is to never store $b$, but rather use $a_{\ell,i}$ to store $b_{\ell,i}$ until hitting line 8: this will work because from that point onwards $b_{\ell,i}$ is never used again and $a_{\ell,i}$ was not needed (or available) thus far. Hence, the static memory overhead is minimal: at least in the LCSM case, one directly aggregates contributions to the same activation tensor, thus not using any extra storage (on top of the inherent $MLD$ it takes to store the activations).
@@ 4 The Flash Inference Framework
Notice that in the case of LCSMs, it never mattered how one computes the contribution of a range of inputs to another range: only that it is well-defined (as soon as the inputs are available) and that it can be done more efficiently than the brute-force approach. Thus, the crux of Algorithm 2 is not convolution-specific, but rather the way we tile the space of contributions, while ensuring autoregressive generation can still take place.
@@ 4.1 Architectural Properties
    Contribution-based The used mixers work by aggregating contributions of each input position to each subsequent output position. That is, for any $\mathrm{mixer}_{\ell}$, there exist an associative aggregation function $\mathrm{agg}:{\mathcal{X}}^{*}\to{\mathcal{X}}$, a read function $\mathrm{read}:{\mathcal{X}}\to\mathbb{R}^{D}$ and a contribution function $cont:\mathbb{R}^{D}\times\mathbb{N}\times\mathbb{N}\to{\mathcal{X}}$ such that:
@@ 4.2 Setting and Definitions
Furthermore, for this choice of ${\mathcal{A}}$, let ${\mathcal{T}}:\mathbb{N}\times\mathbb{N}\to\mathbb{N}$ such that for any $l\leq r<l^{\prime}\leq r^{\prime}$, evaluating $A(y,[l,r],[l^{\prime},r^{\prime}])$ takes at most ${\mathcal{T}}(r-l+1,r^{\prime}-l^{\prime}+1)$ FLOPs.
@@ 4.3 Main Result
Under the assumptions P.1 and P.2, one can generate $L=2^{P}$ tokens autoregressively by performing, per layer, $L-1$ black-box calls to ${\mathcal{A}}$ as well as $L$ more calls to $\mathrm{cont}$, $\mathrm{agg}$, $\mathrm{read}$ and $\mathrm{block}$. Concretely, there are $L/2=2^{P-1}$ calls to ${\mathcal{A}}$ of length $1$ each, $L/4=2^{P-2}$ calls of length $2$ each and so on up to $1$ call of length $L/2$. Hence, neglecting the $\mathrm{cont}$, $\mathrm{agg}$, $\mathrm{read}$ and $\mathrm{block}$ part, the overall time complexity per mixer layer is:
@@
Here, we are being ignorant of the $L$ calls to $\mathrm{cont}$, $\mathrm{agg}$, $\mathrm{read}$ and $\mathrm{block}$ because they only scale linearly with $L$ however MLPs scale quadratically in $D$ and therefore these can be the limiting factor when $D$ is large in comparison to $L$.
The method underlying Theorem 2 is illustrated in Algorithm 4 and follows the same shape as its LCSM counterpart. The only difference is that ${\mathcal{X}}$ is not assumed to be $\mathbb{R}^{D}$ necessarily, so one cannot store the intermediate accumulated states in the same place they store the activations as per the discussion in Section 3.3 (though they could reuse some of the space since $a_{\ell,i}$ and $b_{\ell,i}$ never need to coexist). We use $b_{\ell,i}$ to store the inner part of Equation 6, namely $b_{\ell,i}$ incrementally computes $\mathrm{agg}(\mathrm{cont}(a_{\ell-1},1,i),\mathrm{cont}(a_{\ell-1},2,i),\ldots\mathrm{cont}(a_{\ell-1},i,i))$. Finally, we explicitly moved the gray tile calls to happen after the inner loop and in parallel - this simply shows the last point of the theorem, namely that the calls to ${\mathcal{A}}$ can be done parallelly across layers.
@@
0:  Functions ${\mathcal{A}},\mathrm{cont},\mathrm{agg},\mathrm{read},\mathrm{block}_{[1,M]}$, first token $a_{0,1}$ and $\mathrm{sampler}$
1:  Output: All activations $a_{0},\ldots,a_{M}\in\mathbb{R}^{L\times D}$ obtained by autoregressively sampling
@@
3:  for $i\leftarrow 1$ to $L-1$ do
4:     $U\leftarrow$ maximum power of $2$ that divides $i$ the side of the $i$th gray tile
@@
6:        # account for the contribution of $a_{\ell-1,i}$ to $b_{\ell,i}$ red cell
7:        $b_{\ell,i}=\mathrm{agg}(b_{\ell,i},\mathrm{cont}(a_{\ell-1},i,i))$
8:        $a_{\ell,i}=\mathrm{block}_{\ell}(\mathrm{read}(b_{\ell,i}))$
@@
10:     # account for the contribution of $a_{\ell-1,[i-U+1,i]}$ to $b_{\ell,[i+1,i+U]}$ gray tile
@@ 5 Experiments
We performed two kinds of experiment: one synthetic and one on the Hyena architecture (Poli et al., 2023). The synthetic setup defines all $\mathrm{block}$s to be MLPs with a hidden dimension of $2D$ and GELU activation; it also simply sets $a_{0,i+1}$ as $a_{M,i}$ plus some noise to avoid dependency on vocabulary size since that is out of the scope of our framework and will be negligible at scale. Note that this can be viewed as a sampler: a function from logits at the last layer and previous position to the next token’s embedding. In both settings the weights are initialized to random noise since it does not affect the runtime - this avoid unnecessarily training a new model for each hyperparameter choice.
In terms of notation, we introduce the batch dimension $B$ and keep $M,D,L$ refer to the number of layers, embedding dimension and number of tokens generated, respectively. We use $U$ to refer to the square-tile length, as per line 4. We sweep over $B\in\{1,2,4,8\},M\in\{18,36\}$ [^2] In the Hyena setting, M represents the effective number of mixers, but we fix the Hyena Order to 3 - so M=18 corresponds to 9 Hyena Operators and 18 mixers, $D\in\{256,768\}$ and generate tokens up to the greatest power of $2$ that fits the memory. All results are obtained by averaging over $4$ runs following $2$ runs of warm-up. We evaluate our approach on on the latest NVIDIA H100 and A100 GPUs.

### v2

@@ 3.1.2 Performing LCSM inference via fast relaxed polynomial interpolation
We can now use Algorithm 1 to get an efficient inference algorithm for LCSMs, but to do so we need to be careful about data-dependency. First, for brevity, in our previous notation of activations, disregard the dependency on input $x$ - that is, use $a^{\ell}_{i}$ instead of $a^{\ell}(x)_{i}$ - since we are only going to work with one input $x$ (the one we are generating). Furthermore, denote the intermediate mixer computation by $b\in\mathbb{R}^{M\times L\times D}$ as follows:
$b^{\ell}_{i}\triangleq\mathrm{mixer}^{\ell}(a^{\ell-1})_{i}=\sum_{k=0}^{i}{a^{\ell-1}_{k}\odot\rho^{\ell}_{i-k}}$ (4)
$a^{\ell}_{i}\triangleq\mathrm{block}^{\ell}(b^{\ell}_{i})$ (5)
for any $1\leq\ell\leq M$, where $a^{0}_{i}$ is the embedding of $i$th token. Whenever one has generated the tokens $x_{[1,i]}$ up to position $i$ or, equivalently, $a^{0}_{[1,i]}$, the whole of $a^{[0,M]}_{[1,i-1]}$ will have been computed (since $a^{M}_{[1,i-1]}$ is needed to sample $x_{[1,i-1]}$). Caching these up is the equivalent of a KV-cache in transformers and is simply the natural way to not repeat work when autoregressively sampling. To generate the next token $x_{i+1}=a^{0}_{i+1}$, we want to fill in the values of $a^{[1,M]}_{i}$. To do so, as per Eq. 4, we need to compute $b^{\ell}_{i}$ which is practically, for each level $\ell$, and within each dimension $1\leq c\leq D$, a relaxed interpolation between $a^{\ell-1}_{:,c}$ (playing the role of $y$) and $\rho^{\ell}_{:,c}$. Further applying a $\mathrm{block}^{\ell}$ to $b^{\ell}$, one gets $a^{\ell}$. Thus, we can pipeline several instances of Algorithm 1 sequentially across layers.
The resulting algorithm is illustrated in Algorithm 2. The outer loop (line 3) iterates through positions $i$: at each iteration, one computes the activations at position $i$ across all layers, does some (limited) eager work and samples next token. To do so, the inner loop (line 5) iterates through the several mixer-layers. It first accounts for the red cells in the diagram, that is, the direct dependency on previous layer’s activations at position $i$ (line 7), thus finalizing $b^{\ell}_{i}$. After computing the activation at position $i$ (line 8), it also accounts for the contribution of the gray tile that has just been unlocked (line 10) at the current layer. We reuse the notation from previous section for $\tau$ to be the same as before but within each of the $D$ dimensions - that is:
@@
0: filter $\rho\in\mathbb{R}^{M\times L\times D}$, $\mathrm{block}^{[1,M]}$, first token $a^{0}_{1}$ and $\mathrm{sampler}$
1:Output: All activations $a^{0},\ldots,a^{M}\in\mathbb{R}^{L\times D}$ obtained by autoregressively sampling
@@
3:for $i\leftarrow 1$ to L1L-1 do
4:  $U\leftarrow$ maximum power of $2$ that divides $i$ # the side of the $i$th gray tile
@@
6:   # account for the contribution of $a^{\ell-1}_{i}$ to $b^{\ell}_{i}$ - red cell
7:   $b^{\ell}_{i}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=a^{\ell-1}_{i}\odot\rho^{\ell}_{0}$
8:   $a^{\ell}_{i}=\mathrm{block}^{\ell}(b^{\ell}_{i})$
9:   # account for the contribution of $a^{\ell-1}_{[i-U+1,i]}$ to $b^{\ell}_{[i+1,i+U]}$ - gray tile
10:   $b^{\ell}_{[i+1,i+U]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=\tau(a^{\ell-1},[i-U+1,i],\rho^{\ell},[i+1,i+U])$
@@
The overall time complexity of Algorithm 2 is $O(MDL\log^{2}L)$ for the mixer-part plus LMLM $\mathrm{block}$ calls, to generate $L=2^{P}$ tokens. Furthermore, for each $0\leq q\leq P-1$, there are $MD2^{P-1-q}$ calls to $\tau$ on ranges of length $2^{q}$. The overall amount of memory it takes to store the activations is $O(MLD)$ (as is the case for the naive lazy approach) and the peak memory usage does not change the asymptotics.
@@ 3.2 Across-layer Parallelization
One important feature of the proposed method is that it allows for higher parallelization across layers. In particular, the gray lines can be taken out of the inner loop and performed in parallel across all layers at once, just after the loop (Algorithm 3) - this is because all their inputs and outputs are disjoint. However, the red lines need to be performed sequentially since $a^{\ell}_{i}$ is a function of $b^{\ell}_{i}$ which is a function of $a^{\ell-1}_{i}$ for every $\ell$.
@@
0: filter $\rho\in\mathbb{R}^{M\times L\times D}$, $\mathrm{block}^{[1,M]}$, first token $a^{0}_{1}$ and $\mathrm{sampler}$
1:Output: All activations $a^{0},\ldots,a^{M}\in\mathbb{R}^{L\times D}$ obtained by autoregressively sampling
@@
3:for $i\leftarrow 1$ to L1L-1 do
4:  $U\leftarrow$ maximum power of $2$ that divides $i$ # the side of the $i$th gray tile
@@
6:   # account for the contribution of $a^{\ell-1}_{i}$ to $b^{\ell}_{i}$ - red cell
7:   $b^{\ell}_{i}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=a^{\ell-1}_{i}\odot\rho^{\ell}_{0}$
8:   $a^{\ell}_{i}=\mathrm{block}^{\ell}(b^{\ell}_{i})$
@@
10:  # account for the contribution of $a_{\cdot,[i-U+1,i]}$ to $b_{\cdot,[i+1,i+U]}$ - gray tile(s)
@@
12:     $b^{\ell}_{[i+1,i+U]}\mathrel{\raisebox{0.81805pt}{$\scriptstyle+$}}=\tau(a^{\ell-1},[i-U+1,i],\rho^{\ell},[i+1,i+U])$
@@ 3.3 Memory Considerations
As follows from Proposition 1, the overall size of the inputs to (and outputs of) $\tau$ in the proposed algorithm is only $O(MDL\log L)$ - that is, we access on average activation values at $\log L$ positions per iteration as opposed to the naive implementations (such as lazy and eager approaches) that access an average of $\Omega(L)$ positions. This directly translates to data movement improvements by a factor of up to $L/\log L$.
A simple optimization of static memory is to never store $b$, but rather use $a^{\ell}_{i}$ to store $b^{\ell}_{i}$ until hitting line 8: this will work because from that point onwards $b^{\ell}_{i}$ is never used again and $a^{\ell}_{i}$ was not needed (or available) thus far. Hence, the static memory overhead is minimal: at least in the LCSM case, one directly aggregates contributions to the same activation tensor, thus not using any extra storage (on top of the inherent $MLD$ it takes to store the activations).
@@ 4 The Flash Inference Framework
Notice that in the case of LCSMs, it never mattered how one computes the contribution of a range of inputs to another range: only that it is well-defined (as soon as the inputs are available) and that it can be done more efficiently than the brute-force approach. Thus, the crux of Algorithm 2 is not convolution-specific, but rather the way we tile the space of contributions while ensuring autoregressive generation can take place.
@@ 4.1 Architectural Properties
    Contribution-based The used mixers work by aggregating contributions of each input position to each subsequent output position. That is, for any $\mathrm{mixer}^{\ell}$, there exist an associative aggregation function $\mathrm{agg}:{\mathcal{X}}^{*}\to{\mathcal{X}}$, a read function $\mathrm{read}:{\mathcal{X}}\to\mathbb{R}^{D}$ and a contribution function $cont:\mathbb{R}^{D}\times\mathbb{N}\times\mathbb{N}\to{\mathcal{X}}$ such that:
@@ 4.2 Setting and Definitions
Furthermore, for this choice of $\mathcal{A}$}, let $\mathcal{T}$$\mathbb{N}\times\mathbb{N}\to\mathbb{N}$ such that for any $l\leq r<l^{\prime}\leq r^{\prime}$, evaluating $A(y,[l,r],[l^{\prime},r^{\prime}])$ takes at most ${\mathcal{T}}(r-l+1,r^{\prime}-l^{\prime}+1)$ FLOPs.
@@ 4.3 Main Result
Under the assumptions P.1 and P.2, one can generate $L=2^{P}$ tokens autoregressively by performing, per layer, L1L-1 black-box calls to $\mathcal{A}$} as well as $L$ more calls to $\mathrm{cont}$, $\mathrm{agg}$, $\mathrm{read}$ and $\mathrm{block}$. Concretely, there are $L/2=2^{P-1}$ calls to $\mathcal{A}$} of length $1$ each, $L/4=2^{P-2}$ calls of length $2$ each and so on up to $1$ call of length $L/2$. Hence, neglecting the $\mathrm{cont}$, $\mathrm{agg}$, $\mathrm{read}$ and $\mathrm{block}$ part, the overall time complexity per mixer layer is:
@@
Here, we are being ignorant of the $L$ calls to $\mathrm{cont}$, $\mathrm{agg}$, $\mathrm{read}$ and $\mathrm{block}$ because they only scale linearly with $L$ - however MLPs scale quadratically in $D$ and therefore these can be the limiting factor when $D$ is large in comparison to $L$.
Proof.
Figure 1 shows why the tiling used in Algorithm 2 covers every pair of contributions exactly once and in the correct order (remember, we only assumed $\mathrm{agg}$ to be associative, not necessarily commutative). The more general algorithm is illustrated in Algorithm 4 and follows the same shape as its LCSM counterpart. The only difference is that $\mathcal{X}$} is not assumed to be $\mathbb{R}^{D}$ necessarily, so one cannot store the intermediate accumulated states in the same place they store the activations (though they could reuse some of the space since $a^{\ell}_{i}$ and $b^{\ell}_{i}$ never need to coexist). We use $b^{\ell}_{i}$ to store the inner part of Equation 6, namely $b^{\ell}_{i}$ incrementally computes $\mathrm{agg}(\mathrm{cont}(a^{\ell-1},1,i),\mathrm{cont}(a^{\ell-1},2,i),\ldots\mathrm{cont}(a^{\ell-1},i,i))$. Finally, we explicitly moved the gray tile calls to happen after the inner loop and in parallel - this simply shows the last point of the theorem, namely that the calls to $\mathcal{A}$} can be done parallelly across layers.
The only part remaining to be proved is the right counting of calls to $\mathcal{A}$} per layer: calls of length $2^{q}$ happen whenever $2^{q}$ divides $i$ but $2^{r+1}$ does not - that is, for $i\in\{2^{q},3\cdot 2^{q},5\cdot r^{2},\ldots(2^{P-q}-1)\cdot 2^{q}\}$, so $2^{P-q-1}$ calls. ∎
@@
0: Functions ${\mathcal{A}},\mathrm{cont},\mathrm{agg},\mathrm{read},\mathrm{block}^{[1,M]}$, first token $a^{0}_{1}$ and $\mathrm{sampler}$
1:Output: All activations $a^{0},\ldots,a^{M}\in\mathbb{R}^{L\times D}$ obtained by autoregressively sampling
@@
3:for $i\leftarrow 1$ to L1L-1 do
4:  $U\leftarrow$ maximum power of $2$ that divides $i$ # the side of the $i$th gray tile
@@
6:   # account for the contribution of $a^{\ell-1}_{i}$ to $b^{\ell}_{i}$ - red cell
7:   $b^{\ell}_{i}=\mathrm{agg}(b^{\ell}_{i},\mathrm{cont}(a^{\ell-1},i,i))$
8:   $a^{\ell}_{i}=\mathrm{block}^{\ell}(\mathrm{read}(b^{\ell}_{i}))$
@@
10:  # account for the contribution of $a^{:}_{[i-U+1,i]}$ to $b^{:}_{[i+1,i+U]} -$ gray tile(s)
@@ 5 Experiments
We performed two kinds of experiment: one synthetic and one on the Hyena architecture (Poli et al., 2023). The synthetic setup defines all $\mathrm{block}$s to be MLPs with a hidden dimension of $2D$ and GELU activation; it also simply sets $a^{0}_{i+1}$ as $a^{M}_{i}$ plus some noise to avoid dependency on vocabulary size since that is out of the scope of our framework and will be negligible at scale. Note that this can be viewed as a sampler: a function from logits at the last layer and previous position to the next token’s embedding. In both settings the weights are initialized to random noise since it does not affect the runtime - this avoid unnecessarily training a new model for each hyperparameter choice.
In terms of notation, we introduce the batch dimension $B$ and keep $M,D,L$ refer to the number of layers, embedding dimension and number of tokens generated, respectively. We use $U$ to refer to the square-tile length, as per line 4. We sweep over $B\in\{1,2,4,8\},M\in\{18,36\}$, $D\in\{256,768\}$ and generate tokens up to the greatest power of $2$ that fits the memory. All results are obtained by averaging over $4$ runs following $2$ runs of warm-up. We evaluate our approach on on the latest NVIDIA H100 and A100 GPUs. (Results for all the parameter sweeps in supplementary material).